Calculate a 90% confidence interval for the average age at which gifted children first count to 10 successfully. Choose the closest answer.



1
Expert's answer
2021-10-08T11:04:46-0400
Solution:

Mean=30.69

n=36

SD="\sigma=4.31"

SE"=\sigma/\sqrt n=4.31/\sqrt{36}=0.718"

90% CI value of z is 1.645.

So, z=1.645

So, CI"=(\mu\pm z\times SE)=(30.69\pm1.645\times0.718)"

Thus, CI"=(29.51,31.87)"